OpenStudy (anonymous):

Solve the equation x2 = 81 by finding square roots. For any solution that involves two answers, type one answer separated by a comma and then the other answer

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

do you know the square root of 81?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

you have to remember \[\sqrt{x^2}=\pm x\]

OpenStudy (anonymous):

nope i never did pre-algebra

OpenStudy (anonymous):

it means \[\sqrt{81}= \pm 9\] +9, -9

OpenStudy (anonymous):

2 answers

OpenStudy (anonymous):

wich one should i put tho?

OpenStudy (anonymous):

reread the question

OpenStudy (anonymous):

ok i did

OpenStudy (anonymous):

For any solution that involves two answers, type one answer separated by a comma and then the other answer

OpenStudy (anonymous):

oh right